I have been trying to figure out two problems dealing with a system of three linear equations and I keep getting confused with the steps. I used the elimination and addition method but I'm stuck.

Problem 1 is:                                    Problem 2 is:
                             2x+y-z=10
3x+4y+z=14                 6x+5y-2z=23
  2y+7z=25                   3x-y+z=5
 -2y+3z=5                                                  

Answer by Edwin McCravy(20086)

Problem 1 is:                                    Problem 2 is:
                                                 2x+y-z=10
3x+4y+z=14                                       6x+5y-2z=23
 2y+7z=25                                        3x-y+z=5
-2y+3z=5

Since x is eliminated from both the 2nd and 3rd equation, just solve them:

 2y+7z=25                                        
-2y+3z=5

Add them and the y's cancel and you get

   10z=30
     z=3

Substitute z=3 in

   2y+7z=25
 2y+7(3)=25
   2y+21=25
      2y=4
       y=2


Substitute y=2 and z=3 in 

     3x+4y+z=14
 3x+4(2)+(3)=14
      3x+8+3=14
       3x+11=14
          3x=3
           x=1

(x,y,z) = (1,2,3)

--------------------------------

 Problem 2 is:
2x+ y- z=10
6x+5y-2z=23
3x- y+ z= 5

Add the 1st and 3rd as they are and both the y's and will 
cancel:

2x+y-z=10
3x-y+z= 5
-----------
5x    =15
     x=3

Substitute x=3 in 

  2x+y-z=10
2(3)+y-z=10
   6+y-z=10
     y-z=4 

Substitute x=3 in the middle original equation:

  6x+5y-2z=23
6(3)+5y-2z=23
  18+5y-2z=23
     5y-2z=5

Now you have this system:

      y- z=4
     5y-2z=5

Eliminate z by multiplying the first by -2 and adding:

    -2y+2z=-8
     5y-2z= 5
    ---------
     3y   =-3
         y=-1

Substitute y=-1 in       
       
        y-z=4
     (-1)-z=4
       -1-z=4
         -z=5
          z=-5

(x,y,z) = (3,-1,-5)

Edwin